OpenStudy (anonymous):

square root 2x-1-square root x+2 =0

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

\[\sqrt{2x-1}=\sqrt{x+2}\]

OpenStudy (anonymous):

Then square both sides to remove the roots.

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

So you'll have: 2x - 1 = x + 2 Take care.

OpenStudy (anonymous):

Thank u

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

Don't forget to mark a "Best Response". Thanks!